Circular No. 8355

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2002 T7 (LINEAR) M. L. Sitko, University of Cincinnati; R. W. Russell and D. L. Kim,

Aerospace Corporation; and S. M. Brafford, University of Dayton, report 3–13-μm spectrophotometry of comet C/2002 T7, obtained on May 31.1 UT with the 1.5-m University of Minnesota telescope (+ Aerospace Broadband Array Spectrograph System; 8′′.5 aperture; 49′′ chop throw; integration times 10 min on the comet and 20 min on the reference star, β Gem) at Mt. Lemmon: “A virtually featureless comet continuum was seen to rise from 3.5 to 13 μm, with only a very weak possible silicate emission band superimposed. A blackbody with a temperature of ∼ 300 ± 10 K was fit to the underlying continuum flux at 5, 8.4, and 12 μm. This grain temperature is ∼ 9 ± 4 percent higher than that of an equilibrium blackbody at the heliocentric distance of the comet. The silicate-feature-to-continuum ratio, based on the underlying continuum fit, was ∼ 1.03 ± 0.02. With our aperture, the comet has the following narrowband (∼ 0.25 μm) magnitudes and combined random errors (due to calibration star and comet, as well as variations due to the presence of real spectral structure): [3.7 μm] = 7.63 _{−0.31}^{+0.44}, [4.7 μm] = 5.58 ± 0.09, [5 μm] = 4.79 _{−0.33}^{+0.49}, [8 μm] = 1.43 ± 0.16, [10.5 μm] = 0.06 ± 0.32, and [12 μm] = −0.39 ± 0.02 (the stated errors are standard deviations of the mean).”

COMET C/2004 H6 (SWAN)

Improved parabolic orbital elements from MPC 52169:

T = 2004 May 12.680 TT  ω = 269.^{o}030
Ω = 317.239 }2000.0
q = 0.77605 AU  i = 107.653
2004TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  Mag.
June 14 2^{h}49.^{m}46 −13^{o}25.^{′}6 1.156 0.991 53.^{o}8 55.^{o}8  7.8
19 2 37.14 −13 22.0 1.079 1.050 60.1 57.0  7.8
24 2 22.76 −13 16.4 0.998 1.112 67.0 57.3  7.8
29 2 05.48 −13 08.9 0.916 1.176 74.8 56.5  7.8
July 4 1 44.24 −12 58.9 0.835 1.243 83.7 54.4  7.8

Visual total-magnitude estimates: May 27.80 UT, 7.2 (C. E. Drescher, Warrill View, Qld., 25×100 binoculars); 29.35, 7.3 (A. Amorim, Florianopolis, Brazil, 20×80 binoculars); June 14.37, 7.5 (Amorim).

2004 June 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green